# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

### v1

@@ Abstract
Federated Learning (FL) enables model training across decentralized devices without sharing raw data, thereby preserving privacy in sensitive domains like healthcare. In this paper, we evaluate Kolmogorov–Arnold Networks (KAN) architectures against traditional MLP across six state-of-the-art FL algorithms on a blood cell classification dataset. Notably, our experiments demonstrate that KAN can effectively replace MLP in federated environments, achieving superior performance with simpler architectures. Furthermore, we analyze the impact of key hyperparameters—grid size and network architecture—on KAN performance under varying degrees of Non-IID data distribution. Additionally, our ablation studies reveal that optimizing KAN width while maintaining minimal depth yields the best performance in federated settings. As a result, these findings establish KAN as a promising alternative for privacy-preserving medical imaging applications in distributed healthcare. To the best of our knowledge, this is the first comprehensive benchmark of KAN in FL settings for medical imaging task.
@@ I Introduction
The Kolmogorov–Arnold Networks (KAN) [10, 11], recently proposed as a potential alternative to Multi-Layer Perceptrons (MLP) [12], can theoretically represent continuous functions without using traditional activation functions like other neural networks. While recent studies have used KAN for various centralized uses [13, 14], research on using them in federated settings is still new, mostly limited to the vanilla FedAvg method or using simple datasets [15, 16, 17]. Current research lacks thorough comparisons of KAN over recent FL methods, which prevents a deeper understanding of their practical use in privacy-sensitive medical settings. Addressing this gap is crucial to fully understand KAN’s potential in complex, spread-out medical data environments. Therefore, comprehensive testing over recent FL methods is essential.
To evaluate KAN’s applicability to FL, we compared their performance against MLP-based models across various cutting-edge FL methods, as shown in Fig.  1. This comparison shows that KAN models, even in their simplest form, perform as well as deeper MLP networks when used with various FL methods in IID setting. Additionally, we examined KAN architectures from multiple perspectives, investigating the impact of key hyperparameters—specifically grid size, number of hidden layers—when integrated within federated frameworks in Non-IID nature. These findings will encourage further exploration of KAN’s wider use in FL, especially for complex medical imaging tasks.
@@
  • We demonstrate that KAN compare favorably with MLP across six state-of-the-art FL algorithms.
@@
  • Additionally, we examine whether increasing KAN’s depth through additional hidden layers or expanding its width has greater impact.
@@ II-A Federated Setting
where $F$ denotes the global objective aggregated at the server, and each local objective $f_{n}$ on device $n$ is given by
@@ II-B Federated Learning Process
For clarity and concreteness, we describe the FL framework using the FedAvg algorithm. At each global round $g_{r}\in\{1,2,\dots,G\}$, the central server randomly selects a subset $\mathcal{S}_{g_{r}}\subseteq\{1,2,\dots,N\}$ of edge devices for participation and broadcasts the current global model parameters $\theta_{g}^{(g_{r}-1)}$ to these selected devices.
@@ 4       Server broadcasts $\theta_{g}^{(g_{r}-1)}$ to all devices in $\mathcal{S}_{g_{r}}$
7             Initialize local model: $\theta_{n}^{(g_{r},0)}\leftarrow\theta_{g}^{(g_{r}-1)}$
@@
9            for local epoch $e=0,2,\ldots,E-1$ do
@@ 13                   Compute local loss:
14                   $f_{n}(\theta_{n}^{(g_{r},e)})=\frac{1}{|\mathcal{B}|}\sum_{(x_{i},y_{i})\in\mathcal{B}}\mathcal{L}(h(x_{i};\theta_{n}^{(g_{r},e)}),y_{i})$
@@ 15                   Update local model:
16                   $\theta_{n}^{(g_{r},e+1)}\leftarrow\theta_{n}^{(g_{r},e)}-\eta\nabla f_{n}(\theta_{n}^{(g_{r},e)})$
@@ 24      Server aggregates parameters:
Algorithm 1 FL with Kolmogorov-Arnold Networks
@@ III-A Experiment Setting
To validate the effectiveness of KAN in federated settings, we conduct experiments on the blood cell classification dataset [22] using both KAN and MLP architectures with varying depths. Note that we employed radial basis function-based FastKAN [23] for our experiments to improve computational efficiency. We examine six state-of-the-art FL algorithms: FedAvg [4], FedDyn [24], FedSAM [25], FedGamma [26], FedSMOO [27], and FedSpeed [28] to comprehensively evaluate performance across different federated optimization techniques. Additionally, we simulate a Non-IID data distribution using Dirichlet distribution with $\alpha=1.0$ across $N=100$ edge devices as [29], randomly selecting $10\%$ of devices per global round. We report the mean and standard deviation of test accuracy across five random seeds, and additional hyperparameter settings are available in our open source repository [^1] https://github.com/yjlee22/fkan.
@@ III-C Impact of Non-IID Degree
We investigate how varying the Non-IID degree (Dirichlet parameter $\alpha$) influences the performance of KAN architectures by altering the grid size, as shown in Fig. 3. Results indicate that increasing data heterogeneity (lower $\alpha$) negatively affects test accuracy across all federated learning methods. Larger grid sizes (e.g., $g=10$) generally yield better performance, suggesting grid size is a critical hyperparameter for mitigating Non-IID data challenges. Notably, the improvement from larger grids becomes more pronounced as data heterogeneity intensifies, highlighting grid size’s importance in highly heterogeneous scenarios. Moreover, specific FL algorithms, such as FedGamma and FedSpeed, exhibit notable synergy with larger grid configurations. These findings underscore that appropriately selecting the grid size is essential for optimizing KAN performance in diverse federated learning environments.
@@ III-D Ablation Study
We perform an ablation study using vanilla FedAvg to determine the relative impact of increasing KAN’s depth ($d$) versus expanding its width ($\omega$) as Table. I. As depicted in Fig. 4, results demonstrate that increasing network depth does not guarantee performance improvements, as shallower KAN architectures (1 layer) actually outperform deeper variants (3 and 5 layers) in our experiments. In contrast, width shows significant impact on performance, with the medium configuration (125 parameters) achieving optimal test accuracy compared to both narrow (5 parameters) and wide (3125 parameters) variants. These results highlight the importance of carefully tuning width parameters while keeping depth minimal for enhanced KAN performance in FL scenarios. Note that depth may have a more significant positive impact than observed in our experiments.
@@ IV Conclusion
In this paper, we present the first comprehensive benchmark of KAN in FL for medical image classification. Our results show that KAN can outperform MLP-based network across multiple FL algorithms while maintaining better stability in Non-IID settings. Additionally, we find larger grid configurations provide greater benefits as data heterogeneity increases, offering valuable insights for real-world deployment. Contrary to conventional design principles, our studies reveal that increasing KAN width yields more benefits than increasing depth for medical imaging tasks. Therefore, these findings establish KAN as a promising alternative for privacy-sensitive medical applications, opening new research directions for federated healthcare environments.

### v2

@@ Abstract
Federated Learning (FL) enables model training across decentralized devices without sharing raw data, thereby preserving privacy in sensitive domains like healthcare. In this paper, we evaluate Kolmogorov–Arnold Networks (KAN) architectures against traditional MLP across six state-of-the-art FL algorithms on a blood cell classification dataset. Notably, our experiments demonstrate that KAN can effectively replace MLP in federated environments, achieving superior performance with simpler architectures. Furthermore, we analyze the impact of key hyperparameters—grid size and network architecture—on KAN performance under varying degrees of Non-IID data distribution. In addition, our ablation studies reveal that optimizing KAN width while maintaining minimal depth yields the best performance in federated settings. As a result, these findings establish KAN as a promising alternative for privacy-preserving medical imaging applications in distributed healthcare. To the best of our knowledge, this is the first comprehensive benchmark of KAN in FL settings for medical imaging task.
@@ I Introduction
The Kolmogorov–Arnold Networks (KAN) [10, 11], recently proposed as a potential alternative to Multi-Layer Perceptrons (MLP) [12], can theoretically represent continuous functions without using traditional activation functions like other neural networks. While recent studies have used KAN for various centralized uses [13, 14], research on using them in federated settings is still new, mostly limited to the vanilla FedAvg method or using simple datasets [15, 16, 17]. Current research lacks thorough comparisons of KAN over recent FL methods, which prevents a deeper understanding of their practical use in privacy-sensitive medical settings. Addressing this gap is crucial to fully understand KAN’s potential in complex, spread-out medical data environments. Thus, comprehensive testing of KAN across recent FL methods is essential.
To evaluate KAN’s applicability to FL, we compare their performance against MLP-based models across various cutting-edge FL methods, as shown in Fig. 1. This comparison shows that KAN models, even in their simplest form, perform as well as deeper MLP networks when used with various FL methods in the IID setting. Additionally, we examine KAN architectures from multiple perspectives, investigating the impact of key hyperparameters—specifically grid size and number of hidden layers—when integrated within FL methods in the Non-IID setting. These findings will encourage further exploration of KAN’s wider use in FL, especially for medical imaging tasks requiring AI interpretability.
@@
  • We demonstrate that KAN compare favorably with MLP across six state-of-the-art FL methods.
@@
  • In addition, we examine whether increasing KAN’s depth through additional hidden layers or expanding its width has greater impact.
@@ II-A Federated Setting
where $F$ denotes the global objective aggregated at the server, and each local objective $f_{n}$ on device $n$ is given by:
@@ II-B Federated Learning Process
For clarity and concreteness, we describe the FL framework using the FedAvg method. At each global round $g_{r}\in\{1,2,\dots,G\}$, the central server randomly selects a subset $\mathcal{S}_{g_{r}}\subseteq\{1,2,\dots,N\}$ of edge devices for participation and broadcasts the current global model parameters $\theta_{g}^{(g_{r}-1)}$ to these selected devices.
@@ 4   Server broadcasts $\theta_{g}^{(g_{r}-1)}$ to all devices in $\mathcal{S}_{g_{r}}$
7      Initialize local model: $\theta_{n}^{(g_{r},1)}\leftarrow\theta_{g}^{(g_{r}-1)}$
@@
9    for local epoch $e=1,2,\ldots,E$ do
@@ 13         Compute local loss:
14       $f_{n}(\theta_{n}^{(g_{r},e)})=\frac{1}{|\mathcal{B}|}\!\sum_{(x_{i},y_{i})\in\mathcal{B}}\!\mathcal{L}\!\left(h(x_{i};\theta_{n}^{(g_{r},e)}),y_{i}\right)$
@@ 15         Update local model:
16       $\theta_{n}^{(g_{r},e)}\leftarrow\theta_{n}^{(g_{r},e)}-\eta\,\nabla f_{n}(\theta_{n}^{(g_{r},e)})$
@@ 24  Server aggregates parameters:
Algorithm 1 FL with Kolmogorov–Arnold Networks
@@ III-A Experiment Setting
To validate the effectiveness of KAN in federated settings, we conduct experiments on the blood cell classification dataset [22] using both KAN and MLP architectures with varying depths. Note that we employed radial basis function-based FastKAN [23] for our experiments to improve computational efficiency. We examine six state-of-the-art FL algorithms: FedAvg [4], FedDyn [24], FedSAM [25], FedGamma [26], FedSMOO [27], and FedSpeed [28] to comprehensively evaluate performance across different federated optimization techniques. In addition, we simulate a Non-IID data distribution using Dirichlet distribution with $\alpha=1.0$ across $N=100$ edge devices as [29], randomly selecting $10\%$ of devices per global round. We report the mean and standard deviation of test accuracy across five random seeds, and additional hyperparameter settings are available in our open source repository [^1] https://github.com/yjlee22/fkan.
@@ III-C Impact of Non-IID Degree
We investigate how varying the Non-IID degree (Dirichlet parameter $\alpha$) influences the performance of KAN architectures by altering the grid size, as shown in Fig. 3. Results indicate that increasing data heterogeneity (lower $\alpha$) negatively affects test accuracy across all FL methods. In detail, larger grid sizes (e.g., $g=10$) generally yield better performance, suggesting grid size is a critical hyperparameter for mitigating Non-IID data challenges. Notably, the improvement from larger grids becomes more pronounced as data heterogeneity intensifies, highlighting grid size’s importance in highly heterogeneous scenarios. Moreover, specific FL methods, such as FedGamma and FedSpeed, shows notable synergy with larger grid configurations. These findings underscore that appropriately selecting the grid size is essential for optimizing KAN performance in diverse federated learning environments.
@@ III-D Ablation Study
We perform an ablation study using vanilla FedAvg to determine the relative impact of increasing KAN’s depth ($d$) versus expanding its width ($\omega$) as Table. I. As depicted in Fig. 4, results demonstrate that increasing network depth does not guarantee performance improvements, as shallower KAN architectures (1 layer) actually outperform deeper variants (3 and 5 layers) in our experiments. In contrast, width shows significant impact on performance, with the medium configuration (125 parameters) achieving optimal test accuracy compared to both narrow (5 parameters) and wide (3125 parameters) variants. The results indicate the importance of carefully tuning width parameters while keeping depth minimal for enhanced KAN performance in FL scenarios. Note that depth may have a more significant positive impact than observed in our experiments.
@@ IV Conclusion
In this work, we present the first comprehensive benchmark of KAN in FL for medical image classification. Our results show that KAN can outperform MLP-based network across multiple FL algorithms while maintaining better stability in Non-IID settings. Additionally, we find larger grid configurations provide greater benefits as data heterogeneity increases, offering valuable insights for real-world deployment. Contrary to conventional design principles, our studies reveal that increasing KAN width yields more benefits than increasing depth for medical imaging tasks. Therefore, these findings establish KAN as a promising alternative for privacy-sensitive medical applications, opening new research directions for federated healthcare environments.